# Is it Really Negative? Evaluating Natural Language Video Localization Performance on Multiple Reliable Videos Pool

### v2

@@ 5 MVMR Dataset Analysis
This section describes the used NLVL datasets and detailed settings to derive MVMR positive and negative candidates using SimCSE and EMScore.
NLVL Datasets.
@@
We extend widely-used NLVL datasets (Charades-STA Gao et al. (2017), ActivityNet Krishna et al. (2017), and TACoS Regneri et al. (2013)) to construct three MVMR evaluation datasets, and name the constructed datasets as $\text{MVMR}_{\textsc{\{\$datasource\}}}$. Each NLVL dataset consists of multiple videos with query-moment pairs. The details of each NLVL dataset are shown in Appendix A.
@@
We use STS16 dataset Agirre et al. (2016) to empirically validate varying SimCSE filtering hyper-parameters, $t^{+}_{te}$ and $t^{-}_{te}$. The dataset provides text pairs and human-annotated similarity scores (five-point Likert scale); two texts are regarded as similar when the score is higher than 3. We analyze the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ to derive MVMR positive candidates since it is regarded as the best threshold to distinguish similar queries. We also select $t^{-}_{te}=0.5$ to filter MVMR negative candidates. The distribution of SimCSE and human-annotated scores are shown in Appendix B.
Figure 3: Qualitative Analysis for Filtering Methods. We visualize the derived similarity scores of SimCSE and EMScore to verify the constructed MVMR datasets. We use T-SNE to reduce the dimension of each query embedding for displaying each query (dot) of SimCSE Similarity Distribution.
5.2 MVMR Analysis
We set a video retrieval pool as $n$, containing at most $k$ positive videos for each query. If the number of MVMR positive candidates derived for each query is less than $k$, we include all positive candidates as positive samples. Also, if the total number of MVMR positive and MVMR negative candidates for a query is less than $n$, the query is excluded from the final MVMR dataset.
@@ 5.2 MVMR Analysis
We set the number of videos in the $\text{MVMR}_{\text{Charades-STA}}$ and the $\text{MVMR}_{\text{ActivityNet}}$ retrieval pools as $n=50$, and the maximum number of positive moments per query $k$ is set to $5$. The former consists of 3,716 queries for 1,334 videos, each containing an average of 3.07 positive moments. The latter contains 16,941 queries for 4,885 videos, each including an average of 1.11 positive videos. We set the number of videos for the $\text{MVMR}_{\text{TACoS}}$ and the maximum number of positive moments per query as $5$ and $5$, respectively, since TACoS test set contains only 25 videos. It includes 2,055 queries for 25 videos, each containing an average of 2.24 positive videos. The summary of the MVMR datasets is shown in Table 1.
@@
[…]
Figure 4: Cohesiveness of the Positive and Negative sets. The X and Y axes mean the word co-occurrence ratio and its frequency, respectively. The closer the word’s co-occurrence ratio approaches 1.0, the more cohesive it is.
@@
We recruit crowd workers and ask them to examine 100 queries manually and their paired videos (by our method) and answer whether they are correctly labeled. Overall, we analyze 4,900, 4,900, and 400 videos for Charades-STA, ActivityNet, and TACoS, respectively, since each MVMR dataset consists of 49, 49, and 4 newly added videos (either positive or negative samples). The original video (i.e., golden positive sample) is excluded from this investigation. From human evaluation, we confirm that the newly introduced dataset contains only 1.5%, 0.2%, and 3.5% falsely categorized videos for the three datasets, respectively. TACoS exhibits a relatively higher query-relevant video ratio than the other two datasets since its test set includes only 25 videos and consists of considerably similar cooking activities within the kitchen.
Analysis on Semantic Filtering Methods.
We visualize the scores derived using SimCSE and EMScore to analyze our MVMR datasets qualitatively. Figure 3-(a) displays SimCSE embeddings of all queries in each dataset, illustrating the similarity between all queries and a specific target query, "person turn a light on." (left) and "He is using a push mower to mow the grass." (right). The figure of TACoS can also be found in Appendix B. The figures show that queries exhibit well-defined clusters for Charades-STA and TACoS, revealing the effectiveness of our query similarity-based filtering. These results are also supported by Figure 3-(b) SimCSE score histogram, showing that Charades-STA and TACoS have many similar queries. Figure 3-(c) illustrates histograms of EMScore similarity between queries and videos. The red and blue histograms correspond to the EMScore distribution of positive and negative samples, respectively. For the ActivityNet, EMScore effectively distinguishes positives and negatives. This success is attributed to that ActivityNet includes detailed queries and videos with diverse features. According to the findings presented in this section, the filtering method’s efficacy varies depending on the dataset’s characteristics. Consequently, a combination of two filtering methods (i.g., SimCSE and EMScore) should be employed in a complementary manner to ensure the construction of a reliable MVMR dataset.
6 Methods: CroCs
@@
Figure 5: CroCs Overview. We adopt informative negative contrastive learning to solve the MVMR task. The dots and triangles are the features of moments and texts. The blue dash lines are matched moment-text pairs to be pulled in, while the red dash lines are negative samples of intra/inter-video to be pushed away. The yellow and orange dash lines are unmatched moment-text pairs, but not to train by filtering out since they are easy and false-negatives.
@@
This section introduces a novel contrastive learning method, CroCs, which stands for Cross-directional Hard and Reliable Negative Contrastive Learning, to solve the MVMR task. A key challenge in solving the MVMR task is distinguishing positives from multiple negative distractors. Therefore, we select MMN Wang et al. (2022) since it is the optimal baseline model for the MVMR task as it uses contrastive learning to distinguish negative samples from positives effectively. We use the MMN as the backbone of our model and enhance its ability by applying CroCs in training procedures to solve the challenge in the MVMR task. MMN has limitations in neglecting the potential presence of false-negative video moments since it regards all randomly selected in-batch samples as negatives in contrastive learning. Furthermore, it overlooks the importance gap among negative samples. Our CroCs solves these two limitations of the MMN by adopting two contrastive learning procedures: (1) Weakly-supervised potential negative learning and (2) Informative hard-negative learning.
6.1 Weakly-supervised Contrastive Learning
@@
A critical consideration in the contrastive learning framework is the potential presence of false-negatives. This issue arises when all videos, except for the golden video, are categorized as negative training samples. To mitigate this concern, we propose excluding false-negative videos to refine the contrastive learning process. Specifically, we adopt the SimCSE filtering to identify and exclude false-negative videos as follows:
$V^{-}_{tr}(q)=\{v|s_{te}(q,v)<t_{tr}\},\>\>\ \forall{v}\in V\>$ (1)
@@
where $V^{-}_{tr}(q)$ is a negative sample set of a specific query $q$ to use during the contrastive learning procedure. $s_{te}(q,v)$ is query-video similarity described in Section 4.1. $t_{tr}$ is a hyper-parameter for SimCSE score filtering. If a specific video shows a value of $s_{te}(q,v)$ above the threshold $t_{tr}$ for target query $q$, we classify $v$ as false-negative for $q$. Therefore, we exclude $v$ in negative samples of contrastive learning for $q$.
6.2 Hard-negatives Contrastive Learning
@@
Easy or false-negatives may distract a model from learning proper knowledge during the training process. Inspired by recent progress in using hard-negative sampling for the document retrieval task Zhou et al. (2022), we categorize negatives according to the matching score between a target query and negative moments as follows: (1) Negatives that are clearly irrelevant and have low matching scores should be sampled less frequently (Easy-negative); (2) Negatives that are highly relevant and have high matching scores should also be sampled less frequently (False-negative); (3) Negatives that are uncertain and have matching scores similar to true-positives should be sampled more frequently since they provide useful information (Hard-negative). Based on these criteria, we distinguish hard-negative samples and adopt a retraining process for a model already trained on a specific dataset using only hard-negative samples. We define the negative moments and query sampling distribution as follows:
$p^{\prime}(m|q)\propto\exp{(-(r(q,m)-\overline{r}(q,m^{+}))^{2})},\>\ \forall m\in M^{-}_{q}$ (2)
$p^{\prime}(q|m)\propto\exp{(-(r(q,m)-\overline{r}(q^{+},m))^{2})},\>\ \forall q\in Q^{-}_{m}\>\>$
@@
where $r(q,m)$ is a matching score between a query and a moment calculated using the already trained video moment retrieval model, and $\overline{r}(q,m^{+})$ and $\overline{r}(q^{+},m)$ is mean matching scores for all golden positive moments and queries, respectively. $M^{-}_{q}$ and $Q^{-}_{m}$ is potential negative candidates for a query $q$ and a moment $m$, respectively. $p\textquoteright(m|q)$ means the probability that moment $m$ is sampled as negative when given query $q$. Therefore, negative candidate moments with a close distance to the average matching score of golden query-moment pairs have a high sampling probability. Likewise, $p\textquoteright(q|m)$ means the probability that query $q$ is sampled as negative when given moment $m$. We also reformulate $p\textquoteright(m|q)$ to $p\textquoteright(v|q)$ for computational efficiency in a sampling procedure as follows:
$r(q,v)=\max{(\{r(q,m_{1}),...,r(q,m_{l})\})},\>\>\ \forall m_{i}\in v\>\>\>\>\>\>$ (3)
$p^{\prime}(v|q)\propto\exp{(-(r(q,v)-\overline{r}(q,v^{+}))^{2})},\>\>\ \forall v\in V^{-}_{q}\>\>\>\>\>\>$
@@
$p^{\prime}(q|v)$ can be formulated similarly by calculating $\overline{r}$ about positive queries for each video.
6.3 Reliable Mutual Matching Network
@@
This section describes a model architecture to apply the proposed contrastive learning method, CroCs. We adopt a bi-encoder architecture with a late modality fusion by an inner product in the joint visual-text representational space following the baseline MMN. By adopting a bi-encoder architecture, we can efficiently search for moments with pre-calculated video moment representations given a user query. Our whole model architecture is shown in Figure 5. Our model represents each query and video moment using query and video encoders to two joint visual-text spaces (IoU and mutual matching spaces). From the joint visual-text space, we compute the matching score between the query and video moment representations and select the best video moment. The details of the model architecture are described in the appendix C.1.
@@
We select the binary cross entropy and contrastive learning losses and train our model following two steps: (1) the first training step using the two loss functions (binary cross entropy and contrastive learning losses) filtering out false-negatives described in Section 6.1; (2) the second training step using two loss functions with only hard-negatives described in Section 6.2. To calculate the similarity between the text and video moment features projected in the IoU space, we compute the cosine similarity $s^{iou}$. To adjust the range of the final prediction, we multiply the cosine similarity by a factor of 10, following the MMN. This amplification results in the final prediction $p^{iou}_{i}=\sigma(10\cdot s^{iou}_{i})$, where $\sigma$ represents the sigmoid function. The binary cross-entropy loss is then calculated as follows:
$L_{bce}=-\frac{1}{C}\sum^{C}_{i=1}{(y_{i}\log{p^{iou}_{i}}+(1-y_{i})\log{(1-p^{iou}_{i})})}$ (4)
@@
where $p^{iou}_{i}$ is the confidence score of each moment, and $C$ is the total number of valid candidates. We calculate $L_{bce}$ for positive and true-negative samples derived in Section 6.1.
@@
We use the cross-directional contrastive learning loss to effectively train the model using the informative negatives derived from the methods described in Section 6.1 and 6.2. The mutual matching contrastive learning loss considering the informative negative samples selection process is as follows:
$p^{*}(q_{i}|m)=\frac{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})}{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})+\sum_{j\in\mathcal{I}_{m}^{-}}{\exp(f_{j}^{q\mathsf{T}}f^{m}/\tau_{m})}}\>\>\>$ (5)
$p^{*}(m_{i}|q)=\frac{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})}{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})+\sum_{j\in\mathcal{I}_{q}^{-}}{\exp(f_{j}^{m\mathsf{T}}f^{q}/\tau_{q})}}\>\>\>\>\>$
$L_{rmm}=-(\sum^{N}_{i=1}{(\log{p^{*}(i_{m}|q_{i})}})+\sum^{N}_{i=1}{(\log{p^{*}(i_{q}|m_{i})}}))\>\>\>\>\>\>\>\>\>\>\>$
Dataset Model R@1 R@1 R@1 R@5 R@5 R@5 R@20 R@20 R@20 R@50 R@50 R@50
IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7 IOU0.3 IOU0.5 IOU0.7
$\text{MVMR}_{\text{Charades-STA}}$ CET 3.96 2.96 2.02 15.34 11.79 6.97 39.85 32.21 19.91 59.58 50.54 33.88
BET 4.57 2.99 1.94 14.26 10.98 7.00 31.05 25.11 16.55 45.88 37.86 26.08
G2L 4.30 2.78 1.97 14.34 10.59 7.36 40.89 34.74 21.33 71.92 62.72 44.82
MMN 13.99 12.78 8.40 37.86 33.29 22.36 64.80 57.83 41.71 80.11 73.09 56.75
${}^{\dagger}$ 18.38 16.31 10.76 43.86 39.05 26.99 69.48 62.89 46.39 82.88 76.08 60.28
CroCs 19.27 17.38 11.44 43.54 39.02 27.91 68.08 60.98 46.39 82.99 76.29 60.76
$\text{MVMR}_{\text{ActivityNet}}$ CET 1.26 0.90 0.56 2.81 2.32 1.75 5.41 4.73 3.87 8.97 8.02 6.56
BET 1.56 1.16 0.76 3.02 2.55 2.03 5.30 4.64 3.80 8.13 7.10 5.88
G2L 1.66 1.32 0.96 7.82 6.43 4.55 26.84 21.66 14.81 62.99 50.07 32.99
MMN 13.82 10.71 6.93 31.54 25.04 16.55 54.79 44.50 30.62 73.08 62.09 45.76
${}^{\dagger}$ 16.46 12.68 7.87 36.03 28.47 18.16 57.91 46.52 31.12 73.82 62.12 44.89
CroCs 20.63 15.58 9.51 44.13 34.70 22.40 67.01 55.58 38.55 79.91 70.02 53.20
$\text{MVMR}_{\text{TACoS}}$ CET 8.95 5.79 3.89 26.18 18.88 12.17 54.84 42.53 23.70 - - -
BET 8.22 5.30 2.97 29.46 14.84 8.32 55.67 39.76 21.51 - - -
G2L 8.72 5.82 4.31 30.61 21.51 12.29 56.52 42.35 22.52 - - -
MMN 10.56 8.61 5.60 36.16 25.84 15.47 58.20 43.89 23.80 - - -
${}^{\dagger}$ 13.43 10.75 6.76 36.45 27.01 15.28 58.30 44.38 25.16 - - -
CroCs 13.24 10.75 7.30 39.22 29.68 16.50 60.05 44.77 24.62 - - -
Table 2: Experimental Results on MVMR datasets. Bolded and under-lined results indicate the $1^{st}$ and $2^{nd}$ best performance, respectively. ${}^{\dagger}$ is the model using only weakly-supervised contrastive learning (§6.1). CroCs uses both of two contrastive learning methods (§6.1 and §6.2). We train all the baselines as three trials and report the averaged MVMR scores.
@@
where $q_{i}$ and $m_{i}$ are corresponding positive query and moment for each moment and query, respectively. $f^{m}$ and $f^{q}$ are the moment and query features in the joint visual-text space. $\tau_{q}$ and $\tau_{m}$ are temperatures. $\mathcal{I}^{-}$ means the indices of informative negatives derived using our sampling methods.
@@
The final loss, denoted as $L$, is formulated as a linear combination of the binary cross-entropy loss and the reliable mutual matching loss. The matching score $s$, for a specific moment given the text query, is obtained by multiplying the iou score $s^{iou}$ with the reliable mutual matching score $s^{rmm}$. The reliable mutual matching score is calculated for the features in the mutual matching space in the same way as the iou score.
$L=L_{bce}+\lambda L_{rmm},\>\>\>\>\ s=s^{iou}\cdot s^{rmm}$ (6)
7 Experiments
7.1 Experimental Settings

### v4

@@ 5. MVMR Dataset Analysis
This section describes the utilized VMR datasets for our MVMR framework and detailed settings to derive MVMR positive and negative candidates using SimCSE and EMScore.
VMR Datasets.
@@
We extend widely-used VMR datasets, Charades-STA (Gao et al., 2017), ActivityNet (Krishna et al., 2017), and TACoS (Regneri et al., 2013), to construct three MVMR benchmarks. We name the constructed MVMR datasets as $\text{MVMR}_{\textsc{\{\$datasource\}}}$.
@@
We use the STS16 dataset (Agirre et al., 2016) to empirically validate varying SimCSE filtering hyper-parameters, $t^{+}_{te}$ and $t^{-}_{te}$. The dataset provides text pairs and human-annotated similarity scores (five-point Likert scale), and two texts are considered similar when the score exceeds 3. We analyze the distribution of SimCSE and human annotated scores for all text pairs in the dataset, and select $t^{+}_{te}=0.9$ to derive MVMR positive candidates since it is regarded as the best threshold to distinguish similar queries. Similarly, we select $t^{-}_{te}=0.5$ to identify MVMR negative candidates. The distribution of SimCSE and human-annotated scores are shown in Appendix A.
Figure 4. Qualitative Analysis for Filtering Methods. We visualize the derived similarity scores of SimCSE and EMScore to verify the constructed MVMR datasets. We use T-SNE to reduce the dimension of each query embedding for displaying each query (dot) of SimCSE Similarity Distribution.
5.2. MVMR Analysis
We fixedly set the size of the massive video pool as $n$, containing at most $k$ positive videos for each query. If the number of MVMR positive candidates derived for each query is less than $k$, we include all positive candidates as positive samples. Also, if the total number of MVMR positive and MVMR negative candidates for a query is less than $n$, the query is excluded from the final MVMR dataset.
@@ 5.2. MVMR Analysis
We determine the number of videos in the $\text{MVMR}_{\text{Charades-STA}}$ and the $\text{MVMR}_{\text{ActivityNet}}$ retrieval pools as $n=50$, and the maximum number of positive moments per query $k$ is set to $5$. The former consists of 3,716 queries for 1,334 videos, each containing an average of 3.07 positive moments. The latter contains 16,941 queries for 4,885 videos, each including an average of 1.11 positive videos. For the $\text{MVMR}_{\text{TACoS}}$ dataset, We set the size of the video pool and the maximum number of positive moments per query as $5$ and $5$, respectively, since the TACoS test set contains only 25 videos totally. It includes 2,055 queries for 25 videos, each containing an average of 2.24 positive videos. The summary of the utilized VMR datasets and the constructed MVMR datasets is shown in Table 1.
@@
Analysis on Semantic Distance Verification.
We visualize the results derived using SimCSE and EMScore to analyze our MVMR datasets qualitatively. Figure 4-(a) displays SimCSE embeddings of all queries in each dataset, illustrating the similarity between all queries and a specific target query, ”person turn a light on.” (left) and ”He is using a push mower to mow the grass.” (right). The figure of TACoS can also be found in Appendix A. The figures show that queries exhibit well-defined clusters for Charades-STA and TACoS, revealing the effectiveness of our Query Distance-based filtering. These results are also supported by Figure 4-(b) SimCSE score histogram, showing that Charades-STA and TACoS have many similar queries. Figure 4-(c) illustrates histograms of EMScore similarity between queries and videos. The red and blue histograms correspond to the EMScore distribution of positive and negative samples, respectively. In the case of ActivityNet, EMScore effectively distinguishes positives and negatives. This success is attributed to that ActivityNet includes detailed queries and videos with diverse features. According to the findings presented in this section, the Semantic Distance Verification method’s efficacy varies depending on the dataset’s characteristics. Consequently, a combination of two filtering methods (i.g., SimCSE and EMScore) should be employed in a complementary manner to ensure the construction of a trustful MVMR dataset.
@@
We recruit crowd workers fluent in English through the university’s online community. We ask them to examine 100 queries manually and their paired videos (by our method) and answer whether they are correctly labeled. Overall, we analyze 4,900, 4,900, and 400 videos for Charades-STA, ActivityNet, and TACoS, respectively, since each MVMR dataset consists of 49, 49, and 4 newly added videos (either positive or negative samples). The original videos (i.e., golden positive samples) are excluded from this investigation. Based on human evaluation, we confirm that the newly introduced dataset contains only 1.5%, 0.2%, and 3.5% falsely categorized videos for the three datasets, respectively. TACoS exhibits a relatively higher query-relevant video ratio than the other datasets since its test set includes only 25 videos and consists of significantly similar cooking activities within the kitchen.
6. Methods: CroCs
@@
Figure 5. CroCs Overview. We adopt the informative sample-weighted mutual matching learning to solve the MVMR task. The dots and triangles are the features of moments and texts. The blue dash line is matched moment-text pairs to be pulled in, while the red dash lines are negative samples of intra/inter-video to be pushed away. The yellow and orange dash lines are unmatched moment-text pairs, but not to train by filtering out since they are easy and false-negatives.
@@
To solve the MVMR task, this section introduces a novel informative sample-weighted learning method, CroCs, which stands for Cross-directional informative sample-weighted Contrastive Learning. A key challenge in solving the MVMR task is distinguishing the representation of positives from multiple negative distractors. Therefore, we select MMN (Wang et al., 2022) as the backbone of our model since it is the optimal baseline for the MVMR task since it uses contrastive learning to distinguish negative samples from positives effectively. On the MMN architecture, we apply CroCs in contrastive learning procedures to solve the challenge in the MVMR task. MMN has limitations in neglecting the potential presence of false-negative video moments since it regards all randomly selected in-batch samples as negatives in contrastive learning. Furthermore, it overlooks the importance gap among negative samples. Our CroCs solves these two limitations of MMN by adopting two training mechanisms: (1) Weakly-supervised potential negative learning and (2) Cross-directional hard-negative learning.
6.1. Weakly-supervised Potential Negative Learning
@@
MMN randomly selects negative samples from the whole samples for contrastive learning. This random sampling process has a significant limitation due to the potential presence of false-negatives since it categorizes all samples, except for the golden one, as negative training samples. To mitigate this concern, we propose excluding false-negative videos to refine the contrastive learning process. Specifically, we adopt a potential negative filtering method to identify and exclude false-negative videos as follows:
(1) $V^{-}_{tr}(q)=\{v|s_{te}(q,v)<t_{tr}\},\>\>\ \forall{v}\in V$
@@
where $V^{-}_{tr}(q)$ is a negative sample set of a specific query $q$ to use during the contrastive learning procedure. $s_{te}(q,v)$ is the query-video similarity described in Section 4.1. $t_{tr}$ is a hyper-parameter for the SimCSE score filtering. If a specific video $v$ shows a value of $s_{te}(q,v)$ above the threshold $t_{tr}$ for a target query $q$, we classify $v$ as false-negative for $q$. Therefore, we exclude $v$ in negative samples of contrastive learning for $q$.
6.2. Cross-directional Hard-negative Learning
@@
Easy or false-negative samples may distract a model from learning proper knowledge during the training procedure. Inspired by recent progress in using hard-negative sampling for the document retrieval task (Zhou et al., 2022), we classify negative samples according to the matching score between a target query and negative moments as follows: (1) Negative samples that are clearly irrelevant and have low matching scores should be sampled less frequently (Easy-negative); (2) Negative samples that are highly relevant and have high matching scores should also be sampled less frequently (False-negative); (3) Negative samples that are uncertain and have matching scores similar to true-positives should be sampled more frequently since they provide useful information (Hard-negative). Based on these criteria, we distinguish informative hard-negative samples and adopt a retraining process for a model pre-trained on a specific dataset, using only hard-negative samples. We define the negative sampling distribution for the moment and query as follows:
(2) $p^{\prime}(m|q)\propto\exp{(-(r(q,m)-\overline{r}(q,m^{+}))^{2})},\>\ \forall m\in M^{-}_{q}$
$p^{\prime}(q|m)\propto\exp{(-(r(q,m)-\overline{r}(q^{+},m))^{2})},\>\ \forall q\in Q^{-}_{m}$
@@
where $r(q,m)$ is a matching score between a query and a moment calculated using a pre-trained video moment retrieval model, and $\overline{r}(q,m^{+})$ and $\overline{r}(q^{+},m)$ is mean matching scores for all golden positive moments and queries, respectively. $M^{-}_{q}$ and $Q^{-}_{m}$ is potential negative candidates for a query $q$ and a moment $m$, respectively. $p’(m|q)$ means the probability that a moment $m$ is sampled as negative when given a query $q$. Specifically, $p’(m|q)$ is computed by the distance between $r(q,m)$ and $\overline{r}(q,m^{+})$; thus, the closer the distance, the more frequently sampled as negative since a closer sample is a hard-negative sample not easy to distinguish by a model. In other words, negative candidate moments near the average matching score of golden query-moment pairs have a high sampling probability. Likewise, $p’(q|m)$ means the probability that a query $q$ is sampled as negative when given a moment $m$. We also reformulate $p’(m|q)$ to $p’(v|q)$ for computational efficiency in a sampling procedure as follows:
(3) $r(q,v)=\max{(\{r(q,m_{1}),...,r(q,m_{l})\})},\>\>\ \forall m_{i}\in v$
$p^{\prime}(v|q)\propto\exp{(-(r(q,v)-\overline{r}(q,v^{+}))^{2})},\>\>\ \forall v\in V^{-}_{q}$
@@
$p^{\prime}(q|v)$ can be formulated similarly by calculating $\overline{r}$ for positive queries about each video.
6.3. Informative Sample-weighted Mutual Matching Learning
This section describes an informative sample-weighted contrastive learning method, CroCs.
Model architecture.
@@ 6.3. Informative Sample-weighted Mutual Matching Learning
We adopt a bi-encoder architecture with a late modality fusion by an inner product in the joint visual-text representational space following the baseline MMN. By adopting a bi-encoder architecture, we can efficiently search for moments with pre-computed video moment representations when given a user query. Our whole model architecture is shown in Figure 5. Our model represents each query and video moment using query and video encoders to two joint visual-text spaces (IoU and mutual matching spaces). From the joint visual-text spaces, we compute the matching score between query and video moment representations and select the best video moment. The details of the model architecture are described in the appendix B.1.
Informative Sample-weighted Mutual Matching.
@@
We select the binary cross entropy and contrastive learning losses and train our model following two steps: (1) the first training step using the two loss functions (binary cross entropy and contrastive learning losses) filtering out potential false-negatives described in Section 6.1; (2) the second training step using two loss functions with only informative hard-negatives described in Section 6.2. To calculate the similarity between the text and video moment representations projected in the IoU space, we compute the cosine similarity $s^{iou}$. To adjust the range of the final prediction, we multiply the cosine similarity by a factor of 10, following MMN. This amplification results in the final prediction $p^{iou}_{i}=\sigma(10\cdot s^{iou}_{i})$, where $\sigma$ represents the sigmoid function. The binary cross-entropy loss is then calculated as follows:
(4) $\mathcal{L}_{bce}=-\frac{1}{C}\sum^{C}_{i=1}{(y_{i}\log{p^{iou}_{i}}+(1-y_{i})\log{(1-p^{iou}_{i})})}$
@@
where $p^{iou}_{i}$ is the confidence score of each moment, and $C$ is the total number of valid candidates. We calculate $\mathcal{L}_{bce}$ for positive and potential true-negative samples derived in Section 6.1.
@@
We use the cross-directional contrastive learning loss to effectively train the model using the informative negatives derived from the methods described in Section 6.1 and 6.2. The reliable mutual matching contrastive learning loss considering the informative negative samples selection process is as follows:
(5) $p^{*}(q_{i}|m)=\frac{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})}{\exp((f_{i}^{q\mathsf{T}}f^{m}-\psi)/\tau_{m})+\sum_{j\in\mathcal{I}_{m}^{-}}{\exp(f_{j}^{q\mathsf{T}}f^{m}/\tau_{m})}}$
$p^{*}(m_{i}|q)=\frac{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})}{\exp((f_{i}^{m\mathsf{T}}f^{q}-\psi)/\tau_{q})+\sum_{j\in\mathcal{I}_{q}^{-}}{\exp(f_{j}^{m\mathsf{T}}f^{q}/\tau_{q})}}$
$L_{rmm}=-(\sum^{N}_{i}{\sum^{M}_{j}\log{p^{*}(m_{ij}|q_{i})}}+\sum^{N}_{i}{\sum^{M}_{j}\log{p^{*}(q_{ij}|m_{i})}})$
@@
where $q_{i}$ and $m_{i}$ are corresponding positive query and moment for each moment and query, respectively. $f^{m}$ and $f^{q}$ are moment and query features in the joint visual-text space. $\tau_{q}$ and $\tau_{m}$ are temperatures. $\mathcal{I}^{-}$ means the indices of informative negatives derived using our sampling methods.
@@
The final loss, denoted as $\mathcal{L}$, is formulated as a linear combination of the binary cross-entropy loss and the reliable mutual matching loss. The matching score $s$, for a specific moment given a query, is obtained by multiplying the iou score $s^{iou}$ with the reliable mutual matching score $s^{rmm}$. The reliable mutual matching score is calculated for features in the mutual matching space in the same way as the iou score.
(6) $\mathcal{L}=\mathcal{L}_{bce}+\lambda\mathcal{L}_{rmm},\>\>\>\>\ s=s^{iou}\cdot s^{rmm}$
7. Experiments
7.1. Experimental Settings